# Why Do Open-Source LLMs Struggle with Data Analysis? A Systematic Empirical Study

### v3

@@ 1 Introduction
Large Language Models (LLMs) have shown promise in automating data analysis, with systems like DS-Agent (Guo et al. 2024), AutoML-Agent (Trirat, Jeong, and Hwang 2024), and Data Interpreter (Hong et al. 2024) demonstrating increasingly sophisticated behaviors. Domain-specific benchmarks such as DSBench (Jing et al. 2024), BLADE (Gu et al. 2024), QRData (Liu et al. 2024), and DiscoveryBench (Majumder et al. 2024) further enable targeted evaluation. Yet, performance on these tasks remains dominated by large-scale, advanced models such as GPT-4 (Hurst et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025), while open-source alternatives, especially smaller models, still struggle in real-world analytical settings.
@@ 6 Related Work
To adapt LLMs to complex reasoning tasks, recent work has explored three main approaches: prompt engineering, supervised fine-tuning (SFT), and reinforcement learning (RL). Prompt-based methods (Zhao et al. 2024; Yao et al. 2023; Wang et al. 2023) improve reasoning performance by reformulating inputs to better elicit the model’s latent capabilities. SFT-based methods (Chen et al. 2023; Zeng et al. 2024; Yin et al. 2024; Qiao et al. 2024; Song et al. 2024) provide task adaptation through labeled data, and have been applied in settings such as S1 (Muennighoff et al. 2025), which introduces "budget forcing" to control reasoning steps, and LIMO (Ye et al. 2025), which demonstrates that complex mathematical abilities can emerge from carefully curated examples. RL-based approaches further align model behavior with human preferences and problem-solving strategies. Reinforcement learning further enhances reasoning capabilities by optimizing multi-step decision-making. For example, DeepSeek-R1 (DeepSeek-AI et al. 2025) leverages Group Relative Policy Optimization (GRPO) to train LLMs on complex reasoning tasks. RAGEN (Wang et al. 2025) introduces a modular RL framework that supports stable multi-turn learning via the StarPo architecture. In addition, ReSearch (Chen et al. 2025) integrates RL with retrieval-augmented generation (RAG), treating search operations as part of the reasoning process and learning when and how to retrieve external information.
@@
Data science is an interdisciplinary field that focuses on extracting valuable insights from various data sources, with data analysis serving as a crucial component. Recently, researchers have proposed several specialized LLMs to enhance automated data analysis capabilities (Sun et al. 2024; Li et al. 2025). Data Copilot (Zhang et al. 2023) is a code-centric data analysis agent that efficiently handles massive data processing and visualization through pre-designed interfaces, specifically tailored to automate financial data analysis tasks while reducing errors and improving efficiency. Data Interpreter (Hong et al. 2024) takes a different path by using hierarchical dependency graphs to represent workflows, enabling automatic task breakdown and code improvements. AutoKaggle (Li et al. 2024) coordinates multiple specialized agents (Planner, Developer, and Reviewer) to handle end-to-end data analysis tasks, while keeping room for human input when needed. To evaluate the ability of these agents, researchers have developed several comprehensive benchmarks, such as InfiAgent-DABench (Hu et al. 2024), DSBench (Jing et al. 2024), DA-code (Huang et al. 2024), and DataSciBench (Zhang et al. 2025) for assessing the effectiveness of LLM-based data analysis solutions.

### v4

@@ 1 Introduction
Large Language Models (LLMs) have shown promise in automating data analysis, with systems like DS-Agent (Guo et al. 2024a), AutoML-Agent (Trirat, Jeong, and Hwang 2024), and Data Interpreter (Hong et al. 2024) demonstrating increasingly sophisticated behaviors. Domain-specific benchmarks such as DSBench (Jing et al. 2024), BLADE (Gu et al. 2024), QRData (Liu et al. 2024), and DiscoveryBench (Majumder et al. 2024) further enable targeted evaluation. Yet, performance on these tasks remains dominated by large-scale, advanced models such as GPT-4 (Hurst et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025), while open-source alternatives, especially smaller models, still struggle in real-world analytical settings.
@@ 6 Related Work
To adapt LLMs to complex reasoning tasks, recent work has explored three main approaches: prompt engineering, supervised fine-tuning (SFT), and reinforcement learning (RL) (Wang et al. 2024a; Xi et al. 2025; Guo et al. 2024b; Wang et al. 2024b; Hu et al. 2025; Wang et al. 2025a). Prompt-based methods (Zhao et al. 2024; Yao et al. 2023; Wang et al. 2023) improve reasoning performance by reformulating inputs to better elicit the model’s latent capabilities. SFT-based methods (Chen et al. 2023; Zeng et al. 2024; Yin et al. 2024; Qiao et al. 2024; Song et al. 2024) provide task adaptation through labeled data, and have been applied in settings such as S1 (Muennighoff et al. 2025), which introduces "budget forcing" to control reasoning steps, and LIMO (Ye et al. 2025), which demonstrates that complex mathematical abilities can emerge from carefully curated examples. RL-based approaches further align model behavior with human preferences and problem-solving strategies. Reinforcement learning further enhances reasoning capabilities by optimizing multi-step decision-making. For example, DeepSeek-R1 (DeepSeek-AI et al. 2025) leverages Group Relative Policy Optimization (GRPO) to train LLMs on complex reasoning tasks. RAGEN (Wang et al. 2025b) introduces a modular RL framework that supports stable multi-turn learning via the StarPo architecture. In addition, ReSearch (Chen et al. 2025) integrates RL with retrieval-augmented generation (RAG), treating search operations as part of the reasoning process and learning when and how to retrieve external information.
@@
Data science is an interdisciplinary field that focuses on extracting valuable insights from various data sources, with data analysis serving as a crucial component (Zhang et al. 2025a; Luo et al. 2025; Gupta et al. 2025; Sun et al. 2025). Recently, researchers have proposed several specialized LLMs to enhance automated data analysis capabilities (Sun et al. 2024; Li et al. 2025). Data Copilot (Zhang et al. 2023) is a code-centric data analysis agent that efficiently handles massive data processing and visualization through pre-designed interfaces, specifically tailored to automate financial data analysis tasks while reducing errors and improving efficiency. Data Interpreter (Hong et al. 2024) takes a different path by using hierarchical dependency graphs to represent workflows, enabling automatic task breakdown and code improvements. AutoKaggle (Li et al. 2024) coordinates multiple specialized agents (Planner, Developer, and Reviewer) to handle end-to-end data analysis tasks, while keeping room for human input when needed. To evaluate the ability of these agents, researchers have developed several comprehensive benchmarks, such as InfiAgent-DABench (Hu et al. 2024), DSBench (Jing et al. 2024), DA-code (Huang et al. 2024), and DataSciBench (Zhang et al. 2025b) for assessing the effectiveness of LLM-based data analysis solutions.
@@ 7 Conclusion and Future Work
Acknowledgement
We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023- 00138), Ningbo Natural Science Foundation (2024J020), Yongjiang Talent Introduction Programme (2021A-156-G), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. This work was supported by Ant Group and Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph.